Saturday, July 7, 2012

Day 41, 7/7/2012 -- Dealing with things not working

We came in today to warm cold stages -- this time, neither of our fridge cycle commands worked overnight. After debugging in the morning by Seth, we started new fridge cycles around noon. Most of Kyle's day was spent analyzing old data while Kevin and I worked on debugging issues in the readout electronics. Kyle, Cynthia, and I are back in the high bay at ~9:30PM to try and start a scan and hopefully get some data tonight.

The gondola had its legs attached and the team is doing some initial shakedown tests.

In addition, Kate returned to Palestine, this time with her 1-year-old daughter in tow. Kate and Kevin plan on working out detector issues in the ~week that Kate will be down here.

Pictures (not many):
https://picasaweb.google.com/104253244018605213307/EBEXInPalestine20120707

No comments:

Post a Comment